# Two Pareto Optimum-based Heuristic Algorithms for Minimizing Tardiness and Late Jobs in the Single Machine Flowshop Problem

### v1

@@
Abstract
Flowshop problems play a prominent role in operations research, and have considerable practical significance. The single-machine flowshop problem is of particular theoretical interest. Until now the problem of minimizing late jobs or job tardiness can only be solved exactly by computationally-intensive methods such as dynamic programming or linear programming.
In this paper we introduce, test, and optimize two new heuristic algorithms for mixed tardiness and late job minimization in single-machine flowshops. The two algorithms both build partial schedules iteratively. Both also retain Pareto optimal solutions at intermediate stages, to take into account both tardiness and late jobs within the partial schedule, as well as the effect of partial completion time on not-yet scheduled jobs.
Both algorithms can be applied to scenarios with hundreds of jobs, with execution times running from less than a second to a few minutes. Although they are slower than dispatch rule-based heuristics, the solutions obtained are far better. We also compare a neural-network solution, which performs poorly.
Keywords:
@@ Abstract
1 Background and motivation
1.1 The flowshop problem: an example of a scheduling problem
@@
An important example of a scheduling problem is the flowshop problem. In general, a flowshop may be described as a system in which several “jobs” require processing on several sequential “machines”. In different circumstances “machines” may refer to various types of processors (factory robots, line workers, editors, computer processors, etc.) that perform specified examinations or modifications of objects (automobiles, products, documents, computer programs, etc.), which are designated as “jobs”. In the flowshop problem, all jobs must pass through all the machines one by one in the same sequence. A machine can only process one job at a time. In order to be processed on machine $m$, the job must first pass through machines $11,\ldots,m-$ 1, and it is not possible for two machines to work on the same job at the same time. In practice, flowshop problems occur in a wide variety of manufacturing industries including fabrics, chemicals, electronics, automotive, ([1],[2],[3]), iron and steel [4], food processing [5] ceramic tile [6] packaging and branding. pharmaceuticals and paper manufacture [7] among others.
@@
1.2 Common variants of the single-machine
@@
In different practical situations the flowshop managers may have different objectives. In some cases, the objective may be to minimize makespan, defined as the total time required to complete all jobs. Alternatively, the flowshop customers may set delivery due dates, and the manager may be primarily interested in meeting the customers’ deadlines.
@@
In this paper, we consider situations where all jobs have due times, and each late job has a fixed penalty for being late, plus an additional penalty that is proportional to the amount of time that it is late. There is no bonus for jobs that are finished early. For example if the fixed penalty is 10 and the lateness penalty rate is 5, then a job that is 3 days has cost $10+3\cdot 5=25$.
2 Methods
2.1 Mathematical specification of flowshop model
2.1.1 Constants and variables
@@
  • $p$: fixed penalty (cost) if a job is late;
  • $q$: Lateness penalty coefficient if a job is late. For example, if job $n$ arrives at time $d_{n}+t$, then the lateness penalty for that job is $qt$ (this is in addition to the fixed penalty $p$).
@@
  • Lateness: $w_{k}\in\ {\mathbb{R}}^{+}$ $w_{k}$ is the lateness of the $k$’th job to finish. Note that this is not the same as the job with index $n$. The lateness is always nonnegative ($N$ real variables).
@@
  • Late indicators: $v_{k}\in\{0,1\}$; $v_{k}=1$ if the $k$’th job to complete is late, and 0 otherwise ($N$ binary variables).
2.1.2 Objective function
In this paper, we will consider the problem of minimizing a penalty that depends both on the number of late jobs and on lateness, as described in Section 1.2. In the mathematical specification, this corresponds to the following objective function:
$\sum^{N}_{k=1}{q\cdot w_{k}+}\sum^{N}_{k=1}{{p\cdot v}_{k}}$ (1)
2.1.3 Constraints
@@
  • Positivity of lateness for job in position $k$ ($N$ inequality constraints):
    ${-w}_{k}\leq 0.$ (4)
    This constraint and the previous constraint ensure the correct value of lateness.
  • Late binary indicators ($N$ inequality constraints):
    $w_{k}{-C_{big}}\cdot v_{k}\leq 0\ \forall k,$ (5)
@@
2.2 Algorithms for solution and approximate solution
@@
2.2.1 Existing heuristic scheduling algorithms
@@
Iterative heuristics add some number of jobs at a time and evaluate possible orders before moving on, with the order to be tried determined based on initial criteria, possibly using one of the above algorithms. Existing methods include the Palmer heuristic which calculates an index for each job and uses that to determine the initial order [14], and the NEH heuristic, which prioritizes jobs with longer executions over possibly multiple machines [15]. Until now it appears that the Palmer heuristic has not been adapted to tardiness minmization. On the other hand, there is a modification of NEH (called "NEHedd") that can be applied to tardiness minimization problems [16]. Like NEH, NEHedd proceeds by iteratively constructing longer and longer job sequences, such that the final constructed sequence is the estimated job order. During the iterative process prescribed by NEHedd, it is often the case that more than one partial sequence meets the critierion for selection, i.e. there are "ties" between equally qualified partial sequences. Various mechanisms can be used to resolve these "ties" [17]. However, previous tie-breaking criteria do not make use of Pareto optimality to take into account the multi-criteria nature of the intermediate problems.
2.2.2 Neural network solution for single-machine problem
@@
The NN used is a shallow network consisting of only 3 layers: an input layer, a hidden layer, and an output layer. These three layers are described as follows:
Figure 1: Neural network layer structure for a single-machine system with 8 jobs.
@@
  • Hidden layer: The hidden layer is fully connected. Initial testing led to $4N$ nodes and a linear activation to have the best performance. In initial testing we tried layers with between $4N$ and $8N$ nodes, and with linear, softmin, softmax, softplus, softsign, hyperbolic tangent, SELU, elu, ReLU, and sigmoid activation functions.
  • Output layer: This layer is fully-connected with $N$ nodes, and a sigmoid activation function that produces values between 0 and 1. To convert the output to a job order an “argsort” is performed on the outputs. For example, if the NN produces the output $[0.3,0.4,0.2]$ then the job order should be the third job followed by the first job and then the second job, while the argsort produces $[2,0,1]$ corresponding to the indices.
2.2.3 Pareto-based heuristics for single-machine problem
In this paper we propose two novel heuristic methods, which we designate as "iterative insertion" and "iterative selection".
The iterative insertion algorithm is similar to the NEH algorithm that is used for makespan minimization. Modifications must be made to accommodate special difficulties posed by the lateness-late time minimization criteria. In particular, since multiple criteria are used to determine “good” solutions at intermediate stages, we use Pareto optimality as a selection criterion to determine multiple candidate solutions.
@@ 2.2.3 Pareto-based heuristics for single-machine problem
  • The second list of ordered lists is formed as follows. Using the single list in the first list of lists (which at this stage contains a single element), insert the next preliminary job at each possible point in the list (at this point, there are two possible insertion points), to create two different lists of two elements. For each of these 2-element lists, evaluate the lateness penalty and the finish time associated with these job orderings, and retain all of the Pareto-optimal orderings.
  • Each subsequent stage of the algorithm resembles the second stage. For stage $m$, the $m$’th element of the preliminary list is selected, and inserted at each possible point in all lists in the list of lists that was created at the $m-1$’th stage. For each of the resulting lists, evaluate the lateness penalty and the finish time associated with the given job ordering, and retain all Pareto-optimal orderings.
  • At the last stage, retain only the lists with minimal lateness penalty.
We may give a specific example of the algorithm as follows. Suppose the preliminary order of jobs is $[4,3,2,1]$. At the first stage, the list of lists is $[[4]]$. At the second stage, the lists $[4,3]$ and $[3,4]$ are evaluated for lateness and finish time. Suppose that both latenesses are 0, and the finish times are equal. Then both ordered lists are retained: so at the end of the second stage, the list of ordered lists is $[[4,3],[3,4]]$. At the third stage, job 2 is inserted at each point within each list of ordered lists from the second stage. This produces a list of 6 ordered lists: $[[2,4,3],[4,2,3],[4,3,2],[2,3,4],[3,2,4],[3,4,2]]$. Of these, select those job orderings which are Pareto optimal with respect to lateness and finish time. We may suppose for example that this reduces the list to 3 ordered lists: $[[2,4,3],[2,3,4],[3,2,4]]$. Finally, we insert job 1 at each position in each of these lists, giving rise to 12 ordered lists. From these 12 ordered lists, we select one that has minimum lateness, since each minimum-lateness solution is equally optimal.
@@
When calculating the penalty and finish time for the different jobs with insertion, it is possible to reduce computation time by intelligently choosing the order of calculation. For example, if job 3 is to be inserted in the list $[2,6,4,5,1]$, then one may first calculate the penalty and finish time when the new job 3 is inserted in the last place, i.e. $[2,6,4,5,1,3]$. Next, insert the new job 3 in the second to last place (i.e. $[2,6,4,5,3,1]$–then one may reuse the calculated finish times for $2,6,4,5$, because these jobs’ starting and finishing times remain unchanged. If the calculation is done this way, the complexity is reduced by a factor of 2.
Similarly, if the variant where the starting list consists of $J$ jobs the finish times and latenesses for the $J!$ orderings may be calculated more efficiently by ordering the calculations properly. For example, consider the case where one wants to calculate the penalties and finish times for all permutations of $[1,2,3,4]$. We may first calculate and store the penalties and finish times for each of the single-element ordered lists, namely $[1]$, $[2]$, $[3]$, and $[4]$. Then for example we may use the result for $[1]$ to calculate the finish times and penalties for the 2-element ordered lists that start with 1, namely $[1,2]$, $[1,3]$, $[1,4]$. Similarly, we may use $[2]$ to calculate the finish times and penalties for the 2-element ordered lists that start with 2, namely $[2,1]$, $[2,3]$, $[2,4]$. We may do the same with $[3]$ and $[4]$. Then we may use the results from $[1,2]$ to calculate results for $[1,2,3]$ and $[1,2,4]$: and so on. This method will reduce calculation by 1/2.
@@
  • It is possible that after several iterations, the set of Pareto optimal candidates may become unmanageably large. We counteract this effect by limiting the number of Pareto optimal solutions retained. This is done by retaining a randomly-selected subset of permutations from the Pareto optimal list for each iteration. We thus introduce a tuneable parameter $P$ correponding to the maximum number of permutations retained. In order to maintain a variety among the sequences that are retained, we sort the Pareto optimal candidates by completion time and then select every $\lfloor L/P\rfloor$’th sequence, where $L$ is the number of candidates.
  • One can also limit the insertion slots to a set number $S$, so that new jobs are only inserted into the last $S$ possible positions in the previously-generated job sequences. In especially large numbers of jobs generated with a similar distribution to the training data, sorting by relative due time will put jobs that can be run later closer to the end, and they will not need to be tried at the beginning, as that would just push the other jobs to be late. Rather than trying those jobs at the beginning of the list, time is saved by just inserting in the specified number of insertion slots from the end.
If both of these modifications are implemented, then at each iteration the number of sequences to be calculated is $PS$, and the length of the sequence is $S$ (since it is not necessary to recalculate the job sequence before the first insertion point). It follows that the algorithm’s execution time will be linear in the number of jobs.
@@
Instead of successively building up schedules by inserting new jobs, an alternative approach is based on adding new jobs to candidate schedules sequentially by a similar trial-and-retain procedure. For example, suppose it is desired to schedule 7 jobs $[1,2,3,4,5,6,7]$, and suppose a preliminary ordering (determined by a simple heuristic, such as greedy) has the jobs in order $[2,1,5,6,3,7,4]$. Then for the first $J=4$ jobs, (namely $[2,1,5,6]$ in this case) we may first exhaustively evaluate the penalties and finish times for all $J!=24$ possible orderings of $[2,1,5,6]$. From this list, we may retain the Pareto optimal orderings (with respect to penalty and finish times), and choose the initial job in each of these Pareto optimal orderings. This will give us a list of single-element orderings. Suppose for example, that 2 and 5 are the initial elements of Pareto optimal orderings of $[2,1,5,6]$. We then evaluate finish times and penalties for all orderings of the first five jobs in the preliminary list (namely 2,1,5,6,3) which begin with either 2 or 5. This requires evaluations for $2J!$ orderings. From these, we retain the Pareto optimal orderings, and select the first 2 elements for all of these Pareto optimal orderings. With reference to our example, suppose $[2,6,5,1,3]$, $[2,1,5,6,3]$ and $[5,2,6,1,3]$ are Pareto optimal. Then we will retain the 2-element ordered lists $[2,6]$, $[2,1]$, and $[5,2]$ as candidates for the first two jobs. At the next stage, we evaluate penalties and finish times for all ordered sequences involving the jobs $[2,1,5,6,3,7]$ that begin with either $[2,6]$, $[2,1]$, or $[5,2]$: this involves calculating $3J!$ penalties and finish times. As before, we retain only the first three elements of all Pareto optimal ordered sequences. Suppose for example that all Pareto optimal sequences at this stage begin with $[2,1,6]$. Then at the next stage, we only consider ordered sequences of length 7 that begin with $[2,1,6]$.
@@
As with interative insertion, it is possible that the number of Pareto optimal candidates will grow too large. In this case we may use the same strategy as before: namely order the Pareto optimal candidates by completion time, and choose $P$ regularly-spaced job sequences from among the candidates. It follows that the iterative selection algorithm also has two tuneable parameters, namely the permutation window and the number of retained sequences.
2.3 Training, testing, and evaluation of heuristic methods
2.3.1 Construction of scheduling scenarios for training and testing

### v2

@@
Abstract.
Flowshop problems play a prominent role in operations research, and have considerable practical significance. The single-machine flowshop problem is of particular theoretical interest. Until now the problem of minimizing late jobs or job tardiness can only be solved exactly by computationally-intensive methods such as dynamic programming or linear programming. In this paper we introduce, test, and optimize two new heuristic algorithms for mixed tardiness and late job minimization in single-machine flowshops. The two algorithms both build partial schedules iteratively. Both also retain Pareto optimal solutions at intermediate stages, to take into account both tardiness and late jobs within the partial schedule, as well as the effect of partial completion time on not-yet scheduled jobs. Both algorithms can potentially be applied to scenarios with hundreds of jobs, with execution times running from less than a second to a few minutes. Although they are slower than dispatch rule-based heuristics, the solutions obtained are far better. We also attempted a neural-network solution which performs poorly, and propose reasons why neural networks may not be a suitable approach.
Key words and phrases:
@@ Abstract.
2010 Mathematics Subject Classification:
Primary: 22E30. Secondary: 58J05
Submitted: April 1, 2021. Revised: Apri 10, 2021. Accepted: April 17, 2021.
Correspondence
1. Background and motivation
1.1. The flowshop problem: an example of a scheduling problem
@@
An important example of a scheduling problem is the flowshop problem. In general, a flowshop may be described as a system in which several “jobs” require processing on several sequential “machines”. In different circumstances “machines” may refer to various types of processors (factory robots, line workers, editors, computer processors, etc.) that perform specified examinations or modifications of objects (automobiles, products, documents, computer programs, etc.), which are designated as “jobs”. In the flowshop problem, all jobs must pass through all the machines one by one in the same sequence. A machine can only process one job at a time. In order to be processed on machine $m$, the job must first pass through machines $11,\ldots,m-$ 1, and it is not possible for two machines to work on the same job at the same time. In practice, flowshop problems occur in a wide variety of manufacturing industries including fabrics, chemicals, electronics, automotive, ([1],[2],[3]), iron and steel [4], food processing [5], ceramic tile [6] , packaging and branding, pharmaceuticals, and paper manufacture [7] among others.
@@
1.2. Common variants of the single-machine
@@
In different practical situations the flowshop managers may have different objectives. In some cases, the objective may be to minimize makespan, defined as the total time required to complete all jobs. Alternatively, and the manager may be primarily interested in meeting the customers’ requested deadlines.
@@
In this paper, we consider situations where all jobs have due times, and each tardy job has a fixed penalty for being tardy, plus an additional penalty that is proportional to the amount of time that it is tardy. There is no bonus for jobs that are finished early. For example if the fixed penalty is 10 and the tardiness penalty rate is 5, then a job that is 3 days tardy has cost $10+3\cdot 5=25$.
2. Methods
2.1. Mathematical specification of flowshop model
2.1.1. Constants and variables
@@
  • $p$: fixed penalty (cost) if a job is tardy;
  • $q$: Tardiness penalty coefficient if a job is tardy. For example, if job $n$ arrives at time $d_{n}+t$, then the tardiness penalty for that job is $qt$ (this is in addition to the fixed penalty $p$).
@@
  • Tardiness: $w_{k}\in\ {\mathbb{R}}^{+}$ $w_{k}$ is the tardiness of the $k$’th job to finish. Note that this is not the same as the job with index $n$. The tardiness is always nonnegative ($N$ real variables).
@@
  • Tardy indicators: $v_{k}\in\{0,1\}$; $v_{k}=1$ if the $k$’th job to complete is tardy, and 0 otherwise ($N$ binary variables).
2.1.2. Objective function
In this paper, we will consider the problem of minimizing a penalty that depends both on the number of tardy jobs and on tardiness, as described in Section 1.2. In the mathematical specification, this corresponds to the following objective function:
$\sum^{N}_{k=1}{q\cdot w_{k}+}\sum^{N}_{k=1}{{p\cdot v}_{k}}$ (2.1)
2.1.3. Constraints
@@
  • Positivity of tardiness for job in position $k$ ($N$ inequality constraints):
    ${-w}_{k}\leq 0.$ (2.4)
    This constraint and the previous constraint ensure the correct value of tardiness.
  • Tardy binary indicators ($N$ inequality constraints):
    $w_{k}{-C_{big}}\cdot v_{k}\leq 0\ \forall k,$ (2.5)
@@
    Note that constraint (4) is actually implied by the requirement that $w_{k}\in\ {\mathbb{R}}^{+}$–we have included it in the constraint list for clarity.
2.2. Algorithms for solution and approximate solution
@@
2.2.1. Existing heuristic scheduling algorithms
@@
Iterative heuristics add some number of jobs at a time and evaluate possible orders before moving on, with the order to be tried determined based on initial criteria, possibly using one of the above algorithms. Existing methods include the Palmer heuristic which calculates an index for each job and uses that to determine the initial order [14], and the NEH heuristic, which prioritizes jobs with longer executions over possibly multiple machines [15]. Until now it appears that the Palmer heuristic has not been adapted to tardiness minmization. On the other hand, there is a modification of NEH (called ”NEHedd”) that can be applied to tardiness minimization problems [16]. Like NEH, NEHedd proceeds by iteratively constructing longer and longer job sequences, such that the final constructed sequence is the estimated job order. During the iterative process prescribed by NEHedd, it is often the case that more than one partial sequence meets the critierion for selection, i.e. there are ”ties” between equally qualified partial sequences. Various mechanisms can be used to resolve these ”ties” [17]. However, previous tie-breaking criteria do not make use of Pareto optimality to take into account the multi-criteria nature of the intermediate problems.
2.2.2. Neural network solution for single-machine problem
@@
The NN used is a shallow network consisting of 3 layers: an input layer, a hidden layer, and an output layer. Figure 1 displays the output of tensorflow’s summary command showing the structure of the neural network, indicating the layers, connections, and input and output sizes.
Figure 1. Neural network layer structure for a single-machine system with 8 jobs. Numbers in the right-hand boxes indicate the sizes of input and output arrays for each layer.
A more complete description of the network is as follows.
@@
  • Hidden layer: The hidden layer is fully connected. The number of nodes and the number of hidden layers were selected based on performance for $N=8$ jobs.
  • Output layer: This layer is fully-connected with $N$ nodes, and a sigmoid activation function that produces values between 0 and 1. The estimated schedule is obtained by ordering the outputs.. For example, if the NN produces the output $[0.3,0.4,0.2]$ then the job order should be the third job (lowest value) followed by the first job and lastly the second job (highest value). In the python code, this ordering is recovered using the command argsort.
2.2.3. Pareto-based heuristics for single-machine problem
In this paper we propose two novel heuristic methods, which we designate as ”iterative insertion” and ”iterative selection”.
The iterative insertion algorithm is similar to the NEH algorithm that is used for makespan minimization. Modifications must be made to accommodate special difficulties posed by the tardiness-tardy time minimization criteria. In particular, since multiple criteria are used to determine “good” solutions at intermediate stages, we use Pareto optimality as a selection criterion to determine multiple candidate solutions.
@@ 2.2.3. Pareto-based heuristics for single-machine problem
  • The second list of ordered lists is formed as follows. Using the single list in the first list of lists (which at this stage contains a single element), insert the next preliminary job at each possible point in the list (at this point, there are two possible insertion points), to create two different lists of two elements. For each of these 2-element lists, evaluate the tardiness penalty and the finish time associated with these job orderings, and retain all of the Pareto-optimal orderings.
  • Each subsequent stage of the algorithm resembles the second stage. For stage $m$, the $m$’th element of the preliminary list is selected, and inserted at each possible point in all lists in the list of lists that was created at the $m-1$’th stage. For each of the resulting lists, evaluate the tardiness penalty and the finish time associated with the given job ordering, and retain all Pareto-optimal orderings.
  • At the last stage, retain only the lists with minimal tardiness penalty.
We may give a specific example of the algorithm as follows. Suppose the preliminary order of jobs is $[4,3,2,1]$. At the first stage, the list of lists is $[[4]]$. At the second stage, the lists $[4,3]$ and $[3,4]$ are evaluated for tardiness and finish time. Suppose that both tardinessses are 0, and the finish times are equal. Then both ordered lists are retained: so at the end of the second stage, the list of ordered lists is $[[4,3],[3,4]]$. At the third stage, job 2 is inserted at each point within each list of ordered lists from the second stage. This produces a list of 6 ordered lists: $[[2,4,3],[4,2,3],[4,3,2],[2,3,4],[3,2,4],[3,4,2]]$. Of these, select those job orderings which are Pareto optimal with respect to tardiness and finish time. We may suppose for example that this reduces the list to 3 ordered lists: $[[2,4,3],[2,3,4],[3,2,4]]$. Finally, we insert job 1 at each position in each of these lists, giving rise to 12 ordered lists. From these 12 ordered lists, we select one that has minimum tardiness, since each minimum-tardiness solution is equally optimal.
@@
When calculating the penalty and finish time for the different jobs with insertion, it is possible to reduce computation time by intelligently choosing the order of calculation. For example, if job 3 is to be inserted in the list $[2,6,4,5,1]$, then one may first calculate the penalty and finish time when the new job 3 is inserted in the last place, i.e. $[2,6,4,5,1,3]$. Next, insert the new job 3 in the second to last place, i.e. $[2,6,4,5,3,1]$. Then one may reuse the calculated finish times for $2,6,4,5$, because these jobs’ starting and finishing times remain unchanged. If the calculation is done this way, the complexity is reduced by a factor of 2.
Similarly, if the variant where the starting list consists of $J$ jobs the finish times and tardinesses for the $J!$ orderings may be calculated more efficiently by ordering the calculations properly. For example, consider the case where one wants to calculate the penalties and finish times for all permutations of $[1,2,3,4]$. We may first calculate and store the penalties and finish times for each of the single-element ordered lists, namely $[1]$, $[2]$, $[3]$, and $[4]$. Then for example we may use the result for $[1]$ to calculate the finish times and penalties for the 2-element ordered lists that start with 1, namely $[1,2]$, $[1,3]$, $[1,4]$. Similarly, we may use $[2]$ to calculate the finish times and penalties for the 2-element ordered lists that start with 2, namely $[2,1]$, $[2,3]$, $[2,4]$. We may do the same with $[3]$ and $[4]$. Then we may use the results from $[1,2]$ to calculate results for $[1,2,3]$ and $[1,2,4]$: and so on. This method will reduce calculation by 1/2.
@@
  • It is possible that after several iterations, the set of Pareto optimal candidates may become unmanageably large. We counteract this effect by limiting the number of Pareto optimal solutions retained. This is done by retaining a representative subset of permutations from the Pareto optimal list for each iteration. We thus introduce a tuneable parameter $P$ correponding to the maximum number of permutations retained. In order to maintain a variety among the sequences that are retained, the Pareto optimal candidates are sorted by completion time, and every $\lfloor L/P\rfloor$’th sequence is retained, where $L$ is the number of candidates.
  • One can also limit the insertion slots to a set number $S$, so that new jobs are only inserted into the last $S$ possible positions in the previously-generated job sequences. In especially large numbers of jobs generated with a similar distribution to the training data, sorting by relative due time will put jobs that can be run later closer to the end, and they will not need to be tried at the beginning, as that would just push the other jobs to be tardy. Rather than trying those jobs at the beginning of the list, time is saved by just inserting in the specified number of insertion slots from the end.
The MILP formulation requires $O(N^{2})$ variables and constraints, making exact solutions impractical for $N>10$ (e.g., $N=50$ yields $5,100$ variables). In contrast, the insertion heuristic reduces this to $O(P\cdot S\cdot N)$. If both of these modifications are implemented in the insertion algorithm, then at each iteration the number of sequences to be calculated is $P\cdot S$, and the length of the sequence is $S$ (since it is not necessary to recalculate the job sequence before the first insertion point). It follows that the algorithm’s execution time will be linear in the number of jobs.
@@
Instead of successively building up schedules by inserting new jobs, an alternative approach is based on appending new jobs to candidate partial schedules sequentially by a trial-and-retain procedure. For example, suppose it is desired to schedule 7 jobs $[1,2,3,4,5,6,7]$, and suppose a preliminary ordering (determined by a simple heuristic, such as greedy) has the jobs in order $[2,1,5,6,3,7,4]$. Then for the first $W=4$ jobs, (namely $[2,1,5,6]$ in this case) we may first exhaustively evaluate the penalties and finish times for all $W!=24$ possible orderings of $[2,1,5,6]$. From this list, we may retain the Pareto optimal orderings (with respect to penalty and finish times), and choose the initial job in each of these Pareto optimal orderings. This will give us a list of single-element orderings. Suppose for example, that 2 and 5 are the initial elements of Pareto optimal orderings of $[2,1,5,6]$. We then evaluate finish times and penalties for all orderings of the first five jobs in the preliminary list (namely 2,1,5,6,3) which begin with either 2 or 5. This requires evaluations for $2W!$ orderings. From these, we retain the Pareto optimal orderings, and select the first 2 elements for all of these Pareto optimal orderings. With reference to our example, suppose $[2,6,5,1,3]$, $[2,1,5,6,3]$ and $[5,2,6,1,3]$ are Pareto optimal. Then we will retain the 2-element ordered lists $[2,6]$, $[2,1]$, and $[5,2]$ as candidates for the first two jobs. At the next stage, we evaluate penalties and finish times for all ordered sequences involving the jobs $[2,1,5,6,3,7]$ that begin with either $[2,6]$, $[2,1]$, or $[5,2]$: this involves calculating $3W!$ penalties and finish times. As before, we retain only the first three elements of all Pareto optimal ordered sequences. Suppose for example that all Pareto optimal sequences at this stage begin with $[2,1,6]$. Then at the next stage, we only consider ordered sequences of length 7 that begin with $[2,1,6]$.
@@
As with interative insertion, it is possible that the number of Pareto optimal candidates will grow too large. In this case we may use the same strategy as before: namely order the Pareto optimal candidates by completion time, and choose $K$ regularly-spaced job sequences from among the candidates. It follows that the iterative selection algorithm also has two tuneable parameters, namely the permutation window and the number of retained sequences, which we denote as $W$ and $K$, respectively.
2.3. Training, testing, and evaluation of heuristic methods
2.3.1. Construction of scheduling scenarios for training and testing